# Deterministic generation of single B centers in hBN by one-to-one conversion from UV centers

### v1

@@ VI Conclusion
We have observed electron-beam induced interconversion between individual UV centers and B centers, appearing as abrupt anticorrelated activation or deactivation events of the point defect emission on simultaneously recorded CL scans. This electron-induced switching between these two widely studied color centers shed new light onto their long debated microscopic structure and their mutual relation, consolidating the recent hypothesis that B centers are vertical carbon dimers. Our results provide a direct explanation for the well-known fact that B centers can only be activated in hBN crystals exhibiting the 4.1 eV spectral signature of the UV centers, and suggests a possible activation scenario, where the electron beam would modify the dimer orientation, which would then be made stable by evacuating the leftover vacancy.
@@
In addition, the ability to observe individual activation events has allowed us to make use of the CL signal as a marker heralding the successful activation of single B centers. We illustrated this by generating an array of eight sites with a strongly sub-Poissonnian distribution of the SPE numbers. The rare and undesired double creations can be fixed by selective photobleaching, a laser-induced process that is also heralded by a visible change of the emission signal. The combination of these two activation and deactivation monitoring techniques allows to generate arrays with only one emitter per site. Our results deepen the knowledge of optically active deep defects in hBN and provides clear pathways towards the deterministic fabrication of top-down nanophotonic devices for applications to optical quantum information.
VII Methods
@@
Our sample consists of a single crystal of high-pressure, high-temperature grown hBN [13] exfoliated on a SiO2(300 nm)/Si substrate. After acetone and isopropanol cleaning, the sample is annealed at C under 200 sccm nitrogen (N2) flow during 1 h and left inside the oven during 3 h for cooling. This technique incentivizes the formation of UV emitters. The B centers are created by 3 kV irradiation and an electron current of nA in a commercial scanning electron microscope in high vacuum conditions ( Pa) at room temperature.
VIII Acknowledgments
@@
This work is supported by the French Agence Nationale de la Recherche (ANR) under references ANR-21-CE47-0004. This work received government funding under France 2030 (QuanTEdu-France) with reference number ANR-22-CMAS-0001. The Authors acknowledge K. Watanabe and T. Taniguchi for providing hBN crystals, and thank J.-M. Chauveau and C. Vilar for fruitful discussions.

### v2

@@ Conclusion
The direct observation of electron-induced switching between two widely studied color centers shed new light onto their long debated microscopic structure and their mutual relation, consolidating the recent hypothesis that B centers are vertical carbon dimers. Our results provide a direct explanation for the well-known fact that B centers can only be activated in hBN crystals exhibiting the 4.1 eV spectral signature of the UV centers, and suggests a possible activation scenario, where the electron beam would modify the dimer orientation, which would then be made stable by evacuating the leftover vacancy.
@@
In addition, the ability to observe individual activation events has allowed us to make use of the CL signal as a marker heralding the successful activation of single B centers. We illustrated this by generating an array of eight sites with a strongly sub-Poissonnian distribution of the SPE numbers. The rare and undesired double creations can be fixed by selective photobleaching, a laser-induced process that is also heralded by a visible change of the emission signal. The combination of these two activation and deactivation monitoring techniques allows to generate arrays with only one emitter per site. Our results deepen the knowledge of optically active deep defects in hBN and provides clear pathways towards the deterministic fabrication of top-down nanophotonic devices for applications to optical quantum information. We expect this discussion to trigger further theoretical and experimental studies that could confirm, refine or modify the proposed conversion scenario.
Methods
@@
Our sample consists of single crystals of high-pressure, high-temperature grown hBN.14 This material contains native carbon impurities.19, 11 The flakes are exfoliated on a SiO2(300 nm)/Si substrate. After acetone and isopropanol cleaning (5 min each), the sample is annealed at C under 200 sccm nitrogen (N2) flow during 1 h and left inside the oven during 3 h for cooling. This technique incentivizes the formation of UV emitters.11 The sample is then cleaned again in acetone and isopropanol (5 min each) in an ultrasonic bath.
B center generation
The B centers are created by 3 kV irradiation and an electron current of nA in a commercial scanning electron microscope in high vacuum conditions ( Pa) at room temperature. The CL scans are pixels, sampling a region of m m. The electron irradiation dose is approximately electrons per pixel at each scan. In the single-emitter creation experiment (Fig. 3), the average dose per emitter created is electrons per emitter. In the single SPE array experiment, (Fig. 4), the average dose per emitter created is electrons per emitter.
Optical characterization
Photoluminescence measurements are performed in a room-temperature confocal microscope using a NA = 0.95 air objective. The emitters are excited using a PicoQuant laser diode emitting at nm, and working in either pulsed or cw regime. The photoluminescence signal is filtered using a 17 nm bandpass fluorescence filter centered at 434 nm, and collected by two avalanche photodiodes (Micro Photon Devices PDM and PDM-UV) in the Hanbury Brown and Twiss configuration.
Intensity autocorrelation is measured for all SPEs to ensure that we only consider individual emitters (i.e. ). The functions are fitted using multi-peak functions of the form
where with the laser period of 13.1 ns, the emitter lifetime, and is the amplitude coefficient for the period . and are fit parameters accounting for the signal and background intensities. All coefficients are fixed to 1 except from the central period coefficient , which is the fit parameter providing the value .
Associated content
The present work is associated with a preprint. A. Núñez Marcos, C. Arnold, J. Barjon, S. Buil, J.-P. Hermier, A. Delteil, Deterministic generation of single B centers in hBN by one-to-one conversion from UV centers, 2025, 2511.02465. arXiv. https://arxiv.org/abs/2511.02465
Data availability
The data generated in this study are available at 10.5281/zenodo.19110631.
Supporting information
Supporting information available: additional CL maps under different irradiation conditions; additional data about controlled photobleaching; time-resolved ensemble CL measurements; SEM image of a generated array.
Acknowledgments
@@
This work is supported by the French Agence Nationale de la Recherche (ANR) under references ANR-21-CE47-0004 and ANR-25-CE47-3104. This work received government funding under France 2030 (QuanTEdu-France) with reference number ANR-22-CMAS-0001. The Authors acknowledge K. Watanabe and T. Taniguchi for providing hBN crystals, and thank J.-M. Chauveau and C. Vilar for fruitful discussions.
Supporting Information
Deterministic generation of single B centers in hBN by one-to-one conversion from UV centers
Andrés Núñez Marcos, Christophe Arnold, Julien Barjon, Stéphanie Buil, Jean-Pierre Hermier, Aymeric Delteil
Université Paris-Saclay, UVSQ, CNRS, GEMaC, 78000, Versailles, France.
——————– aymeric.delteil@usvq.fr——————–
I S1. Additional CL maps
In this section we provide eight additional consecutive CL maps recorded on the 145 nm flake (Figure S1). As in the main text, the circles mark some emitters with the same color code (i.e. green for active emitter, red for inactive, two-colored when the state is switching). The measurement protocol is identical to that of Figure 2 of the main text. A minor difference is that the scanned region drifted noticeably during the measurement, which we did not compensate for.
In addition, on Fig. S2 we show CL maps recorded under a much larger current, up to 1.16 nA. Several switching events between UV and B centers are visible on the same map, with some emitters undergoing multiple switching events during the scan.
Finally, we show two blue-range scans of non-irradiated areas, which show that the sample is initially devoid of B centers (Fig. S3). They appear during the scans, which translates in spots where the upper part is missing. All CL maps are scanned from left to right, then from top to bottom.
II S2. Controlled deactivation by photobleaching
In this section, we further illustrate the principle of controlled photobleaching. Figure S4a shows a confocal map of a pre-irradiated area (with randomly positioned emitters). We selected a few emitters and applied the same protocol as in Section III of the main text. Figure S4b shows six examples of timetraces recorded during the 7.5 mW laser illumination. All show an abrupt bleaching of the signal after a few tens to hundreds of seconds. A confocal map recorded after the selective bleaching is shown on Figure S4c, where the missing emitters are denoted by the dashed circles.
III S3. Individual vs ensemble CL measurements
In this section, we compare the CL characterization of emitter population dynamics in the regime where individual emitters can be resolved, with the prior approaches based on ensembles 23, 24, 31, where the implicit assumption is that the CL signal is proportional to the defect population. We first show that the excitation and detection efficiency is in general not constant, before studying how this affects the inferred population kinetics in ensemble measurements.
III.1 Time-dependent excitation and collection efficiency
The efficiency with which an emitter is excited and its light collected during CL measurements typically evolves in time, due to e.g. irradiation-induced surface contamination by carbonaceous adsorbates (see section S4), creation of competitive non-radiative recombination channels, and drift of the electron beam position. Fig. S5 show the amplitude and background associated with single B and UV centers as extracted from Gaussian fit of consecutive CL maps. The intensity of both B and UV centers decrease by 10 to 20 % over the course of the measurement. A confirmation of this time-dependent efficiency is also visible in Fig. 3 of the main text, where the plateaus associated with the single steps exhibit a slight negative slope.
The main consequence of this observation is that the CL signal associated with a given defect species is not purely proportional to the population, but takes the form , where is the CL signal, is the time-dependent efficiency, the population and the background. Note that the latter is also time-dependent.
III.2 Monte-Carlo simulations
In order to qualitatively evaluate the influence of on ensemble measurements, we perform Monte-Carlo simulations of the CL signal. We consider a partially reversible conversion B UV. Fig. S6a and b show the simulated populations as a function of time for three different conversion rates (the simulation parameters are described in the caption). In all cases, we observe time-dependent populations exhibiting anticorrelated fluctuations between the UV and B populations. All curves are well fitted by a monoexponential function, which provides identical time constants for the UV and B population kinetics.
We then plot the associated signal in the realistic hypothesis that decreases by about 10 % during the course of the measurement. Fig. S6c and d show the calculation results. The CL signal is only slightly modified by the introduction of . However, the main consequence is that the results of a monoexponential fit provides vastly different time constant between the UV and B defects, with the UV decay time being found longer than the B decay time (up to a factor 3). This is due to the variations in increasing (decreasing) the curvature of blue (UV) signal. The exponential fits are very sensitive to these slight changes in curvature and yield very different optimal parameters. In the next subsection, we further illustrate this phenomenon with experimental data.
III.3 Ensemble measurements
We perform ensemble CL measurements on a crystal possessing a high density of UV centers. We use a focused beam of current 10 to 20 pA during 200 to 500 s. The results are shown on Fig. S7. As in the simulations, the UV and B signals exhibit opposite time evolutions, with the blue signal increasing while the UV signal decreases. Clear anticorrelations are also visible in the signal fluctuations, with many occurrences of discrete jumps in opposite directions in B and UV signals. However, when fitting the signal with a monoexponential function, different time constant are found for the blue and UV signal, with the timescale associated with the UV signal being consistently longer than that of the blue signal, in agreement with the simulations. This shows a limitation of ensemble measurements when evaluating the kinetics of luminescent species in CL. Our approach, based on the identification of individual emitters using enhanced detection sensitivity associated with spatial resolution, allows access to individual particles, which can then be followed in time, and therefore makes it possible to circumvent the limitations of ensemble measurements.
IV S4. Surface contamination during irradiation
The irradiation process is well known to generate adsorption of carbon-based contaminants on the sample surface. While in-situ monitoring allows to minimize it by stopping the irradiation immediately after the emitter is created, we nevertheless observe a slight contamination at the position of the irradiation sites (Fig. S8).